# An Extensive Study on D2C: Overfitting Remediation in Deep Learning Using a Decentralized Approach

### v1

@@
Abstract
Overfitting remains a significant challenge in deep learning, often arising from data outliers, noise, and limited training data. To address this, we propose Divide2Conquer (D2C), a novel technique to mitigate overfitting. D2C partitions the training data into multiple subsets and trains identical models independently on each subset. To balance model generalization and subset-specific learning, the model parameters are periodically aggregated and averaged during training. This process enables the learning of robust patterns while minimizing the influence of outliers and noise. Empirical evaluations on benchmark datasets across diverse deep-learning tasks demonstrate that D2C significantly enhances generalization performance, particularly with larger datasets. Our analysis includes evaluations of decision boundaries, loss curves, and other performance metrics, highlighting D2C’s effectiveness both as a standalone technique and in combination with other overfitting reduction methods. We further provide a rigorous mathematical justification for D2C’s underlying principles and examine its applicability across multiple domains. Finally, we explore the trade-offs associated with D2C and propose strategies to address them, offering a holistic view of its strengths and limitations. This study establishes D2C as a versatile and effective approach to combating overfitting in deep learning. Our codes are publicly available at: https://github.com/Saiful185/Divide2Conquer.
Index Terms:
Classification, Deep Learning, Hyperparameter, Overfitting, Regularization, Variance.
I Introduction
@@
Deep Learning models often do well on the data they train on, but the performance drops massively on unseen data from different distributions. This phenomenon is known as Overfitting. Many methods have been proposed to reduce overfitting over the years [10]. Early Stopping [27] is one of the most intuitive methods for reducing overfitting. However, it limits the model’s learning potential [28]. Network reduction reduces variance by simplifying the model, but it restricts learning of complex features [29]. Data augmentation is a common approach to mitigate overfitting in deep learning applications [32]. However, choosing suitable augmentation techniques for specific datasets can be challenging, and obtaining additional training data often requires substantial effort. Regularization is a widely used technique to reduce overfitting by penalizing model dependency on training data. Dropout, the most relevant regularization method in deep learning, reduces overfitting by randomly dropping connections between neural network layers [4]. However, even dropout can sometimes degrade the performance of a model due to internal covariance shifts [18]. Eventually, despite employing these techniques, the model still has access to the entire training data, and the robust neural network structures find their ways to fit on the training set a bit too much. This is where our motivation comes from. It can be a good idea to not let the neural network train on the entire dataset in the first place. Intuitively, we should combine multiple models that train on different portions of the data so that the model can not familiarize itself with all of it at once, and a consensus can be achieved by focusing on the general patterns observed in different portions of the data.
@@
We propose a novel method, Divide2Conquer (D2C), which proposes dividing the training data into multiple subsets and training a model(all having the same architecture & hyperparameters) on each subset. A weighted averaging of all the parameter weights from the instance model is performed after a certain number of epochs, and the averaged weights are shared back to each of the models (we denote them as edge models in this study). Then, the whole process is repeated for several global epochs.
D2C is loosely inspired by Federated Optimization[2], which involves training identical models on local devices (using the data available on those devices) and periodically aggregating and averaging their trained parameters on a central server. This process ensures safe access to data and preserves privacy, as it only shares the weights, not the data itself. However, unlike D2C, federated learning is not inherently designed to enhance model generalization.
The D2C method also has similarities with Bagging[23] in the form of using subsets of data. However, in bagging (Bootstrap Aggregating), the data is not completely separated for different models as we proposed. Instead, the technique involves creating multiple subsets of the original dataset by randomly sampling with replacement. This means that some datapoints may be repeated in a subset while others may be excluded altogether. Also, in bagging, the aggregation is done by combining the output probabilities, not the model weights.
@@
Our experiments encompass multiple datasets across different domains, and each case suggests that employing the D2C method can reduce overfitting significantly. The primary comparison to evaluate the method was between the performances of the base Neural Network architecture used for the Edge Models using the entire training data and the performance of our models using the Divide2Conquer method. We summarize our contributions through this study below:
  • We introduce a new method, D2C, that helps in reducing overfitting significantly while being conceptually simple and easy to implement.
  • We present a mathematical justification for the hypothesis that D2C mitigates overfitting by rigorously demonstrating its capacity to reduce model variance. This formal analysis underscores D2C’s potential to improve model robustness across diverse datasets.
@@
  • D2C can be applied on top of other data augmentation and regularization techniques, and our experiments across various datasets and tasks show that this results in a clear improvement in the model’s generalization ability.
  • We extensively tune the hyperparameters introduced by this method and report the findings, providing important directions for future applications.
@@
  • We also delve into the trade-offs associated with D2C, including its computational overhead and potential diminishing returns for smaller datasets. We explore strategies to mitigate these challenges and provide a comprehensive understanding of D2C’s strengths and weaknesses.
@@
The paper is organized as follows. In section II, we discuss relevant literature. We establish the theoretical justification behind our hypothesis in section III. In sections IV and V, we discuss our methodology and lay out the experimental specifications. In section VI, we evaluate our approach through empirical experiments on multiple datasets and analyze the results. We discuss the limitations of D2C in section VII. Section VIII contains concluding remarks.
II Related Works
Several authors have addressed the issue of overfitting while performing various tasks. M. Cogswell et al. proposed a regularizer called DeCov, which helps reduce overfitting in deep neural networks by Decorrelating Representations[5].
Dropout [4] was proposed by Srivastava et al. back in 2014, and since then, this technique has been extensively used in very complex neural network architectures successfully. It was indeed an outstanding contribution, specifically for deep learning-based models. Batch Normalization [8] proposed by Ioffe and Sergei primarily focuses on better convergence and somewhat contributes to reducing overfitting. J. Kolluri and V.K Kotte came up with L¼ regularization to solve the problems faced by L1 and L2 regularization techniques [15].
Ensembles are also often used to improve generalization. The ensemble-based ELVD model [3] managed to outperform the traditional VGGNet and DropoutNet models in terms of reducing overfitting. Zehong Zeng et al. came up with an ensemble framework that incorporates several techniques to prevent overfitting [9]. Min-Gu Kim et al. also proposed parallel ensemble networks to reduce overfitting in ECG data and prevent the degradation of generalization performance as the training progresses [1].
We implemented a method based on federated optimization preliminarily for facial expression recognition using FedNet[17]. This model achieves excellent results in terms of generalization on both CK+ and FER-2013 datasets. The K-Fold-Cross-Validation method for evaluation proposed by Korjus K. et al. [13] has some similarities to our proposed method in the sense that it also creates subsets by partitioning the dataset and makes use of the whole training data. However, this method is different from ours since the erroneous samples still make it to the training phase (K-1/K)-th of the time in K-Fold CV. The proposed D2C method partitions data and the whole training process is partitioned too, which means none of the subsets is fed with a particular outlier while training bar one, minimizing the impact of these erroneous samples.
Overfitting remains a prevalent issue in supervised machine learning despite methods like Early Stopping, Network Reduction, Training Set Expansion, Regularization, and Dropout being effective [10]. Through D2C, we build an approach that focuses on achieving even better generalization and can be implemented on top of other overfitting-reducing techniques.
III Theoretical Framework
III-A The Hypothesis
Overfitting happens when a model fits too well on the data that it trains on. It effectively captures and memorizes even the random characteristics from the training data, randomness that would be insignificant in real-world applications. The presence of outliers and noisy data is a fundamental reason behind overfitting. In this study, we formulate a method to minimize the effect of these outliers and noisy data.
The contribution of outliers/noise can be minimized by dividing the training data into multiple shards and training each shard separately. That way, each data point will only occur once in one of the several data shards. The representative samples would be close to each other in the feature space and would be present in each of the data shards more or less uniformly. However, the individual outliers/noise would only be able to impact one of the training processes. Aggregation and averaging can be done periodically to combine all the models. In the best-case scenario, due to averaging, the effect of the individual outliers/noise would be reduced by a factor of N, where N is the number of data shards. The key factors behind our hypothesis and considerations regarding this method being able to address overfitting are discussed as follows:
@@ III-A The Hypothesis
1. Weighted averaging of parameters helps in combining the knowledge learned from different subsets. However, extreme parameter updates driven by noise or outliers in individual models are moderated when averaging across multiple models. If a model encounters a noisy sample that deviates significantly from the overall data distribution, it might adjust its parameters excessively to fit that sample, leading to overfitting. However, if the training data is divided into subsets and trained separately, even if one model’s parameters are influenced by noisy or outlier samples, their impact is diluted when combined with the parameter weights of other models during averaging. So, the averaged parameters should reflect a more balanced representation of the underlying data distribution, mitigating the influence of individual noisy or outlier samples.
@@
2. Averaging the parameter weights is particularly helpful in combating overfitting since it has a regularizing effect. Averaging weights regularizes the model at the parameter level. It moderates extreme weight updates driven by noise or outliers in individual models, leading to more stable and generalizable parameter values. This helps prevent overfitting by discouraging models from fitting to the noise or idiosyncrasies present in their training subsets. By averaging parameter weights, the central model converges to a shared solution that reflects the collective knowledge learned from different perspectives. This consensus learning approach encourages models to learn generalizable patterns and reduces reliance on individual model predictions that are prone to overfitting.
3. Averaging parameters smooths out the decision boundary learned by individual models. Extreme or noisy parameter updates that result in sharp or jagged decision boundaries in individual models are likely to be moderated and smoothed when combined through averaging, consequently helping in generalizing well to unseen data and reducing overfitting.
4. The optimal hyperparameters, such as the number of subsets and epochs before each round of central averaging vary by dataset and domain. The appropriate subset count depends on dataset size and class distribution. Suppose the edge model is complex or the subset count is too high. In that case, the overly reduced size and number of samples per class in each subset can result in significant overfitting in edge models, which in turn can degrade overall performance even after central averaging. This is similar to what happens in federated learning when the client data is minimal, as pointed out by Zhang et al. [34]. When sufficient data exists per class, more subsets can be beneficial, as noise or outliers are diluted by a factor of up to N (number of training subsets).
In summary, D2C helps to reduce overfitting by mitigating the effects of noise and outliers. However, dividing the training set into too many subsets can lead to insufficient data per subset, which can hamper the overall performance.
III-B Mathematical Intuition
@@
To lay out the underlying intuition behind our hypothesis, we must first define overfitting formally. We follow the groundwork laid out by Ghojogh et al. [23]. Let us consider a Dataset $D$ with training set $T$ and test set $R$ so that,
$T\cup R=D$ (1)
$T\cap R=\emptyset$ (2)
@@
Let, $T$ consist of datapoints $\{(x_{i},y_{i})\}_{i=1}^{n}$ and test set $R$ consist the datapoints $\{(x_{i},y_{i})\}_{i=1}^{m}$. Also, let $f$ be the function that maps each datapoint in the training set to its corresponding output label(true observations), i.e., $f(x_{i})=f_{i}$, where $f_{i}$ is the true corresponding output label for $x_{i}$. This model is unknown and we wish to train a model that replicates $f$ as closely as possible. However, both the ground truth model and the true labels are unknown to us. Realistically, there is always some noise in the training dataset. We assume that the training output is corrupted with some additive noise $\epsilon_{i}$. So, we can say,
$y_{i}=f_{i}+\epsilon_{i}$ (3)
@@
where $\epsilon_{i}$ is some Gaussian noise, $\epsilon\sim\mathcal{N}(0,\sigma^{2})$. Since the mean of the distribution is zero we can say that $\mathop{{}\mathbb{E}}(\epsilon_{i})=0$, and $Var(\epsilon_{i})=\sigma^{2}$. As the variance of a random variable $x$ can be denoted as $Var(X)=\mathop{{}\mathbb{E}}(x^{2})-[\mathop{{}\mathbb{E}}(x)]^{2}$ (Appendix A), we have,
$\mathop{{}\mathbb{E}}(\epsilon_{i}^{2})=Var(\epsilon_{i})+[\mathop{{}\mathbb{E}}(\epsilon_{i})]^{2}\Rightarrow\mathop{{}\mathbb{E}}(\epsilon_{i}^{2})=\sigma^{2}+0=\sigma^{2}$ (4)
@@
Our assumption is that the input of the training dataset ${\{{x}_{i}}\}_{i=1}^{n}$ and their noise-added outputs ${\{{y}_{i}}\}_{i=1}^{n}$ are available to us, and our goal is to estimate the true model $f$ by a model $\hat{f}$ in order to predict the labels ${\{{y}_{i}}\}_{i=1}^{n}$ for the input data ${\{{x}_{i}}\}_{i=1}^{n}$. Let the estimated observations be ${\{\hat{y}_{i}}\}_{i=1}^{n}$. It is preferable that ${\{\hat{y}_{i}}\}_{i=1}^{n}$ are as close as possible to ${\{{y}_{i}}\}_{i=1}^{n}$. Mathematically, we can consider this as a minimization problem looking to minimize some error function such as MSE (Mean squared error). The MSE of estimated outputs from the trained model with respect to the dataset outputs are: $\textbf{MSE}=\frac{1}{n}\sum_{i=1}^{n}(\hat{y}_{i}-y_{i})^{2}=\mathop{{}\mathbb{E}}((\hat{y}_{i}-y_{i})^{2})$. Let us consider a single arbitrary datapoint $(x_{0},y_{0})$ and the corresponding prediction of the trained model $\hat{f}$ for the input $x_{0}$ is $\hat{y}_{0}$. Ghojogh et al. [23] showed that if the sample instance $(x_{0},y_{0})\notin T$ i.e., it belongs to the test set $R$, MSE of the predicted label can be expressed as -
$\textbf{MSE}=\mathop{{}\mathbb{E}}((\hat{y}_{0}-y_{0})^{2})=\mathop{{}\mathbb{E}}((\hat{y}_{0}-f_{0})^{2})+\sigma^{2}$ (5)
@@
Here we can observe from the first term of the R.H.S. of the equation that the MSE of the estimation of output labels with respect to the dataset labels can actually be represented by the MSE of the estimation with respect to the true uncorrupted labels ${\{{f}_{i}}\}_{i=1}^{n}$ plus some effect of the noise that exists in the dataset. Taking Monte-Carlo approximation [24] of the expectation terms in Eq. (5), we have:
@@
$\implies\sum_{i=1}^{m}{(\hat{y}_{i}-y_{i})^{2}}=\sum_{i=1}^{m}{(\hat{y}_{i}-f_{i})^{2}}+m\sigma^{2}$ (6)
Here the term $\sum_{i=1}^{m}{(\hat{y}_{i}-y_{i})^{2}}$ is the total error between the model predictions and the dataset labels which we can call the empirical observed error, $e$. On the other hand, $\sum_{i=1}^{m}{(\hat{y}_{i}-f_{i})^{2}}$ represents the total error between prediction labels and the true unknown labels, namely $E$. Hence, $e=E+m\sigma^{2}$. Thus, the observed error truly reflects the true error because the term $m\sigma^{2}$ remains constant, which is the theoretical justification for evaluating model performance using unseen test data. But, as Ghojogh et al. [23] showed, if $(x_{0},y_{0})\in T$, deriving the M.S.E yields:
$\textbf{MSE}=\mathop{{}\mathbb{E}}((\hat{y}_{0}-y_{0})^{2})=\mathop{{}\mathbb{E}}((\hat{y}_{0}-f_{0})^{2})+\sigma^{2}-2\sigma^{2}\mathop{{}\mathbb{E}}(\frac{\partial\hat{y}_{0}}{\partial y_{0}})$ (7)
@@
$\implies\sum_{i=1}^{n}{(\hat{y}_{i}-y_{i})^{2}}=\sum_{i=1}^{n}{(\hat{y}_{i}-f_{i})^{2}}+n\sigma^{2}-2\sigma^{2}\sum_{i=1}^{n}\frac{\partial\hat{y}_{i}}{\partial y_{i}}$ (8)
@@
Figure 1: When the corresponding complexity term grows and dominates the expression in Eq.(8) for empirical error, the model starts to overfit as empirical error becomes no longer a true representation of true error.
@@
Figure 2: A visualization of (a) Underfitting, (b) A good fit, and (c) Overfitting scenarios, illustrating how the model changes if an arbitrary training data point $(x,y)$ is shifted to $(p,q)$ by plotting the trained model before (solid line) and after (dashed line) the shift.

### v2

# DUA-D2C: Dynamic Uncertainty Aware Method for Overfitting Remediation in Deep Learning

@@
Abstract.
Background: Overfitting remains a significant challenge in deep learning, often arising from data outliers, noise, and limited training data. To address this, we previously proposed the Divide2Conquer (D2C) method, which partitions training data into multiple subsets and trains identical models independently on each. This strategy enables learning more consistent patterns while minimizing the influence of individual outliers and noise.
Objectives: D2C’s standard aggregation typically treats all subset models equally or based on fixed heuristics (like data size), potentially underutilizing information about their varying generalization capabilities. Building upon this foundation, we introduce Dynamic Uncertainty-Aware Divide2Conquer (DUA-D2C), an advanced technique that refines the aggregation process.
Methods: DUA-D2C dynamically weights the contributions of subset models based on their performance on a shared validation set, considering both accuracy and prediction uncertainty. This intelligent aggregation allows the central model to preferentially learn from subsets yielding more generalizable and confident edge models, thereby more effectively combating overfitting.
Results: In this work, we provide a rigorous theoretical justification for DUA-D2C’s efficacy. Empirical evaluations on benchmark datasets demonstrate that DUA-D2C significantly improves generalization, even when applied on top of other regularization methods. Our analysis includes evaluations of decision boundaries, loss curves, and other performance metrics, highlighting the effectiveness of DUA-D2C.
Conclusions: This study demonstrates that DUA-D2C improves generalization performance even when applied on top of other methods, establishing it as a theoretically grounded and effective approach to combating overfitting in deep learning.
copyright: ccdoi: 10.1613/jair.1.xxxxxpublicationmonth: 9journalyear: 2025
1. Introduction
@@
Deep Learning models often achieve high performance on their training data, yet this performance can drop significantly on unseen data from different distributions, a phenomenon known as overfitting. Numerous methods have been proposed over the years to mitigate overfitting (ying2019overview, 30). Early Stopping (morgan1989generalization, 21), while intuitive, may limit a model’s learning potential (prechelt2002early, 24). Network reduction simplifies models to reduce variance but can restrict the learning of complex features (furnkranz1997pruning, 7). Data augmentation is a common approach (sun2014deep, 28), but selecting suitable techniques can be challenging, and acquiring additional data often requires substantial effort. Regularization, particularly methods like Dropout (srivastava2014dropout, 27) which randomly deactivates connections, is widely used. However, even Dropout can sometimes degrade performance due to issues like internal covariance shifts (li2019understanding, 17). A common thread is that despite these techniques, the model ultimately trains on the entire dataset, allowing robust neural networks to potentially over-adapt to the training set.
This observation motivates our work. We posit that preventing the neural network from training on the entire dataset simultaneously can be beneficial. Instead, by combining multiple models trained on different portions of the data, a consensus can be achieved, focusing on general patterns rather than allowing any single model to become overly familiar with the complete dataset.
@@
To this end, we previously proposed the Divide2Conquer (D2C) method (siddiqui2024d2c, 1). D2C partitions the training data into multiple subsets and trains a model (all instances sharing the same architecture and hyperparameters) on each subset. After a certain number of local training epochs, the parameters from these edge models (models trained on subsets) are aggregated through weighted averaging, and the resulting averaged weights are disseminated back to each edge model. This cycle is repeated for several global epochs. In this study, apart from proposing DUA-D2C, we provide a rigorous mathematical justification for the hypothesis that D2C mitigates overfitting by demonstrating its capacity to reduce model variance.
While D2C provides a robust framework for improving generalization, its standard aggregation mechanism typically treats all edge models equally based on their data subset size. To further refine this process and enhance model generalization, we introduce an extension: Dynamic Uncertainty-Aware Divide2Conquer (DUA-D2C). DUA-D2C enhances the aggregation step by dynamically weighting the contributions of edge models. These weights are determined by evaluating each edge model’s performance on a shared validation set, considering both its predictive accuracy and its prediction uncertainty (e.g., measured via entropy). This intelligent, adaptive weighting scheme allows the central model to preferentially learn from edge models that demonstrate superior generalization and confidence, thereby more effectively minimizing the influence of outliers, noise, and subset-specific overfitting.
The D2C approach was inspired by Federated Optimization (konevcny2015federated, 13), which trains models on local devices and aggregates parameters centrally, primarily for privacy. However, unlike D2C and DUA-D2C, federated averaging is not inherently designed as an overfitting reduction technique. D2C also shares similarities with Bagging (ghojogh2019theory, 8) in its use of data subsets. However, Bagging typically involves sampling with replacement, meaning data is not strictly separated, and aggregation is performed on output probabilities rather than model weights.
@@
Our experiments encompass multiple datasets across different domains. Each case suggests that employing the D2C framework, particularly with the DUA-D2C enhancement, can significantly reduce overfitting. The primary comparison to evaluate the methods was between the performance of the base neural network architecture (used for the Edge Models) trained on the entire dataset and the performance of our models utilizing the D2C and DUA-D2C approaches.
We summarize our contributions through this study below:
  • We introduce Dynamic Uncertainty-Aware D2C (DUA-D2C), a novel extension that employs an intelligent, dynamic weighting scheme based on edge model validation performance (accuracy and uncertainty) to further enhance model robustness and generalization.
  • We present a rigorous mathematical justification for the Divide2Conquer approach and extend it to highlight DUA-D2C’s effectiveness, demonstrating its ability to reduce model variance and thereby improve generalization.
@@
  • We demonstrate through extensive experiments that DUA-D2C can be applied on top of other data augmentation and regularization techniques, leading to clear improvements in model generalization across various datasets and tasks.
@@
  • We delve into the trade-offs associated with the DUA-D2C, such as computational overhead and potential diminishing returns for smaller datasets, exploring strategies to mitigate these challenges and offering a comprehensive understanding of their strengths and weaknesses. We also tune the hyperparameters introduced by DUA-D2C and report the findings, providing important directions for future applications.
@@
The paper is organized as follows. We establish the theoretical justification behind the D2C hypothesis in Section 2. In Section 3, we detail the D2C methodology and its DUA-D2C enhancement, followed by the experimental specifications in Section 4. In Section 5, we evaluate our approaches through empirical experiments on multiple datasets and analyze the results. We discuss the limitations of our methods in Section 6. In Section 7, we discuss relevant literature. Section 8 contains concluding remarks.
2. Theoretical Framework
2.1. The Hypothesis
Overfitting occurs when a model adapts excessively to its training data, effectively capturing and memorizing random characteristics and noise that lack significance in real-world applications. The presence of outliers and noisy data is a fundamental contributor to this phenomenon. Our approach aims to minimize the detrimental impact of such data irregularities.
The core idea is that the influence of outliers and noise can be substantially mitigated by partitioning the training data into multiple shards and training distinct models on each. In this setup, each data point is confined to a single shard. While representative samples, being closer in the feature space, are likely to be distributed somewhat uniformly across shards, individual outliers or noisy instances will primarily affect only one of the local training processes. Through periodic aggregation of these locally trained models, we can consolidate their learned knowledge.
For the basic Divide2Conquer (D2C) approach, which employs simple averaging, the adverse effect of individual outliers/noise could ideally be reduced by a factor approaching N (the number of data shards). Building upon this, our Dynamic Uncertainty-Aware D2C (DUA-D2C) introduces a more sophisticated aggregation mechanism. Instead of uniform averaging, DUA-D2C dynamically weights the contribution of each edge model based on its performance (accuracy) and predictive confidence (inverse uncertainty/entropy) on a shared validation set. This refinement aims to further amplify the benefits of the Divide2Conquer paradigm.
The key factors underpinning our hypothesis that the DUA-D2C method addresses overfitting are as follows:
@@ 2.1. The Hypothesis
I. In D2C, weighted averaging of parameters combines knowledge from different subsets. Extreme parameter updates driven by noise or outliers in individual models are moderated during this averaging. If an edge model encounters a noisy sample that deviates significantly, it might adjust its parameters excessively. However, this impact is diluted when combined with parameters from other models. The DUA-D2C extension enhances this by actively down-weighting models that exhibit poor generalization (low accuracy on validation) or high uncertainty (potentially due to fitting noise in their local shard). This ensures that the aggregated parameters more reliably reflect a balanced representation of the underlying data distribution, offering a more robust mitigation against individual noisy or outlier samples.
@@
II. Averaging weights has a regularizing effect at the parameter level. D2C’s averaging moderates extreme weight updates, leading to more stable parameter values. DUA-D2C refines this by making the regularization adaptive. By prioritizing models that are both accurate and confident on unseen validation data, DUA-D2C discourages the central model from adopting parameters heavily influenced by noise or idiosyncrasies from less reliable edge models. This intelligent consensus learning encourages the central model to learn more generalizable patterns and reduces reliance on individual model predictions that might be prone to overfitting their local data.
III. Averaging parameters generally smooths out the decision boundaries learned by individual models. Sharp or jagged boundaries resulting from noisy updates in individual models are moderated through D2C’s averaging. DUA-D2C provides an additional layer of quality control: models that likely developed overly complex or noisy decision boundaries (as indicated by poor validation performance or high uncertainty) contribute less to the central model. This leads to a central model with a potentially smoother and more robust decision boundary, better equipped for generalization to unseen data.
IV. The optimal hyperparameters, such as the number of subsets $N$ and local epochs $E$, remain crucial and vary by dataset and domain. The subset count depends on dataset size and class distribution. If edge models are complex or the subset count is too high, the reduced data per subset can lead to significant overfitting in these edge models. This is similar to what happens in federated learning when the client data is minimal, as pointed out by Zhang et al. (zhang2022fed, 34). Statistically, the variance of an estimated model $Var(\hat{f}_{j})$ is often inversely related to the amount of data used for its estimation. When $n_{j}$ is small, the parameters $\theta_{j}$ learned by $\hat{f}_{j}$ can be highly sensitive to the specific instances in $D_{j}$. Small perturbations in these few samples can lead to large changes in the learned $\theta_{j}$, indicating high variance. The model might learn spurious correlations present only in that small $D_{j}$ or effectively memorize its samples, rather than capturing the underlying, generalizable patterns from the true data distribution $P(x,y)$. This local overfitting manifests as $Var(\hat{f}_{j})$ being large, meaning $\hat{f}_{j}$ would produce widely different predictions if trained on a different small random sample from $P(x,y)$. When sufficient data exists per class ($n_{j}$ is reasonably large), more subsets can be beneficial, as noise or outliers are diluted by a factor of up to N (number of training subsets), and each $\hat{f}_{j}$ can learn more stable representations. While D2C’s averaging helps, DUA-D2C offers a mechanism to better handle this scenario. If some edge models overfit due to limited data in their shard, their likely poor performance or high uncertainty on the validation set will reduce their influence on the central model, offering a degree of resilience against poorly performing edge models. Nevertheless, ensuring sufficient data per class in each subset remains important for all edge models to learn meaningful representations. The dynamic weighting in DUA-D2C aims to leverage the ”best” of what each edge model learns, even if some are suboptimal.
In summary, DUA-D2C’s dynamic uncertainty-aware aggregation helps reduce overfitting by mitigating the effects of noise and outliers and by promoting a more robust, generalized central model. While dividing the training set into too many subsets can still pose challenges, DUA-D2C provides a more sophisticated mechanism to capitalize on the contributions of potentially variably performing edge models.
2.2. Mathematical Intuition
@@
To lay out the underlying intuition behind our hypothesis, we must first define overfitting formally. We follow the groundwork laid out by Ghojogh et al. (ghojogh2019theory, 8). Let us consider a Dataset $D$ with training set $T$ and test set $R$ so that,
(1) $T\cup R=D$
(2) $T\cap R=\emptyset$
@@
Let, $T$ consist of datapoints $\{(x_{i},y_{i})\}_{i=1}^{n}$ and test set $R$ consist the datapoints $\{(x_{i},y_{i})\}_{i=1}^{m}$. Also, let $f$ be the function that maps each datapoint in the training set to its corresponding output label(true observations), i.e., $f(x_{i})=f_{i}$, where $f_{i}$ is the true corresponding output label for $x_{i}$. This model is unknown, and we wish to train a model that replicates $f$ as closely as possible. However, both the ground truth model and the true labels are unknown to us. Realistically, there is always some noise in the training dataset. We assume that the training output is corrupted with some additive noise $\epsilon_{i}$. So, we can say,
(3) $y_{i}=f_{i}+\epsilon_{i}$
@@
where $\epsilon_{i}$ is some Gaussian noise, $\epsilon\sim\mathcal{N}(0,\sigma^{2})$. Since the mean of the distribution is zero we can say that $\mathop{{}\mathbb{E}}(\epsilon_{i})=0$, and $Var(\epsilon_{i})=\sigma^{2}$. As the variance of a random variable $x$ can be denoted as $Var(X)=\mathop{{}\mathbb{E}}(x^{2})-[\mathop{{}\mathbb{E}}(x)]^{2}$ (see Appendix A), we have,
(4) $\mathop{{}\mathbb{E}}(\epsilon_{i}^{2})=Var(\epsilon_{i})+[\mathop{{}\mathbb{E}}(\epsilon_{i})]^{2}\Rightarrow\mathop{{}\mathbb{E}}(\epsilon_{i}^{2})=\sigma^{2}+0=\sigma^{2}$
@@
Our assumption is that the input of the training dataset ${\{{x}_{i}}\}_{i=1}^{n}$ and their noise-added outputs ${\{{y}_{i}}\}_{i=1}^{n}$ are available to us, and our goal is to estimate the true model $f$ by a model $\hat{f}$ in order to predict the labels ${\{{y}_{i}}\}_{i=1}^{n}$ for the input data ${\{{x}_{i}}\}_{i=1}^{n}$. Let the estimated observations be ${\{\hat{y}_{i}}\}_{i=1}^{n}$. It is preferable that ${\{\hat{y}_{i}}\}_{i=1}^{n}$ are as close as possible to ${\{{y}_{i}}\}_{i=1}^{n}$. Mathematically, we can consider this as a minimization problem looking to minimize some error function, such as MSE (Mean Squared Error). The MSE of estimated outputs from the trained model with respect to the dataset outputs are: $\textbf{MSE}=\frac{1}{n}\sum_{i=1}^{n}(\hat{y}_{i}-y_{i})^{2}=\mathop{{}\mathbb{E}}((\hat{y}_{i}-y_{i})^{2})$. Consider a single arbitrary data point $(x_{0},y_{0})$ and the corresponding prediction of the trained model $\hat{f}$ for the input $x_{0}$ is $\hat{y}_{0}$. Ghojogh et al. (ghojogh2019theory, 8) showed that if the sample instance $(x_{0},y_{0})\notin T$ i.e., it belongs to the test set $R$, MSE of the predicted label can be expressed as -
(5) $\textbf{MSE}=\mathop{{}\mathbb{E}}((\hat{y}_{0}-y_{0})^{2})=\mathop{{}\mathbb{E}}((\hat{y}_{0}-f_{0})^{2})+\sigma^{2}$
@@
Here we can observe from the first term of the R.H.S. of the equation that the MSE of the estimation of output labels with respect to the dataset labels can actually be represented by the MSE of the estimation with respect to the true uncorrupted labels ${\{{f}_{i}}\}_{i=1}^{n}$ plus some effect of the noise that exists in the dataset. Taking Monte-Carlo approximation (ghojogh2020sampling, 9) of the expectation terms in Eq. (5), we have:
@@
(6) $\implies\sum_{i=1}^{m}{(\hat{y}_{i}-y_{i})^{2}}=\sum_{i=1}^{m}{(\hat{y}_{i}-f_{i})^{2}}+m\sigma^{2}$
Here the term $\sum_{i=1}^{m}{(\hat{y}_{i}-y_{i})^{2}}$ is the total error between the model predictions and the dataset labels which we can call the empirical observed error, $e$. On the other hand, $\sum_{i=1}^{m}{(\hat{y}_{i}-f_{i})^{2}}$ represents the total error between prediction labels and the true unknown labels, namely $E$. Hence, $e=E+m\sigma^{2}$. Thus, the observed error truly reflects the true error because the term $m\sigma^{2}$ remains constant, which is the theoretical justification for evaluating model performance using unseen test data. But, as Ghojogh et al. (ghojogh2019theory, 8) showed, if $(x_{0},y_{0})\in T$, deriving the M.S.E yields:
(7) $\textbf{MSE}=\mathop{{}\mathbb{E}}((\hat{y}_{0}-y_{0})^{2})=\mathop{{}\mathbb{E}}((\hat{y}_{0}-f_{0})^{2})+\sigma^{2}-2\sigma^{2}\mathop{{}\mathbb{E}}(\frac{\partial\hat{y}_{0}}{\partial y_{0}})$
@@
(8) $\implies\sum_{i=1}^{n}{(\hat{y}_{i}-y_{i})^{2}}=\sum_{i=1}^{n}{(\hat{y}_{i}-f_{i})^{2}}+n\sigma^{2}-2\sigma^{2}\sum_{i=1}^{n}\frac{\partial\hat{y}_{i}}{\partial y_{i}}$
@@
Figure 1. When the corresponding complexity term grows and dominates the expression in Eq.(8) for empirical error, the model starts to overfit as empirical error becomes no longer a true representation of true error.
@@
Figure 2. A visualization of (a) Underfitting, (b) A good fit, and (c) Overfitting scenarios, illustrating how the model changes if an arbitrary training data point $(x,y)$ is shifted to $(p,q)$ by plotting the trained model before (solid line) and after (dashed line) the shift.